Find moles of insoluble CH3COOAg for Mixture #1 Initial (Agt] 160mm Initial (CH3C00] 1140 m4 Final [Ag+] comm Final (CH3C00] goma TABLE 6.3 Solubility Analysis ofmixture#1

In first table concentration is in mM and table 2 M.

concentration of KSCN
Answer #1

Agt + CH₃ Cool CH3CooAg W 140 - XX 140-x=80 - x = 140-80 l 60 _- . 60-1=0 x=60 - CH3600.ng = 60mm_= 60X10-ºm -0.06m eqb (Ag